OpenStudy (anonymous):

please help what is the fourth root of 8 in the form x + iy?

OpenStudy (anonymous):

will the calculations differ from -8 /

OpenStudy (math&ing001):

Yes because 4 is an even number

OpenStudy (math&ing001):

I think we should put (x+iy)^4=8 and find x and y

OpenStudy (anonymous):

I think this is wrong ? z⁴ = -8 z⁴ + 8 = 0 (z²)² – (2i√2)² = 0 →→→ Because i²=-1 (z² + 2i√2)(z² – 2i√2) = 0 →→→ a²–b²=(a–b)(a+b) [z² – (1 – 2i – 1)(√2)][z² – (1 + 2i – 1)(√2)] = 0 [z² – (1² – 2i + i²)(√2)][z² – (1² + 2i – i²)(√2)] = 0 [z² – (1 – i)²(∜2)²][z² – (1 + i)²(∜2)²] = 0 →→→ a²±2ab+b²=(a±b)² [z – (1 – i)(∜2)][z + (1 – i)(∜2)][z – (1 + i)(∜2)][z + (1 + i)(∜2)] = 0 →→→ a²–b²=(a–b)(a+b)

OpenStudy (math&ing001):

This is actually correct with a little typo mistake on 6th line [z² – (1² – 2i + i²)(√2)][z² – (1² + 2i + i²)(√2)] = 0 (+ instead of -) But this is for the 4th root of -8 and we're looking for the 4th root of 8.

OpenStudy (math&ing001):

Let me think about it for a moment, I'll let you know if I have an answer !

OpenStudy (radar):

Note how the plot converted the 8 to 2^3.

OpenStudy (math&ing001):

Ok I was able to do it manually here's how : z⁴ = 8 z⁴ - 8 = 0 (z²)² – (2√2)² = 0 (z² + 2√2)(z² – 2√2) = 0 →→→ a²–b²=(a–b)(a+b) [z² – (i√2∜2)²][z² – (√2∜2)²] = 0 [z – (i√2∜2)][z + (i√2∜2)][z – (√2∜2)][z + (√2∜2)] = 0 →→→ a²–b²=(a–b)(a+b)

OpenStudy (anonymous):

that makes sense but last question.... if I have to get the value of z1 = then how will I get it

OpenStudy (math&ing001):

Basic equation solving : [z – (i√2∜2)][z + (i√2∜2)][z – (√2∜2)][z + (√2∜2)] = 0 So you have 4 solutions : z1 = i√2∜2 z2= -i√2∜2 z3= √2∜2 z4 = -√2∜2

OpenStudy (anonymous):

thanks that was a stupid question. thanks for your help.

OpenStudy (math&ing001):

Anytime :)